# Trial Length, Pricing, and Rationally Inattentive Customers

### v2

@@ Appendix C. Proofs for the Paid-Trial Extension
We assume $\eta^{\prime}(P_{0})<0$ and sign-up elasticity $\varepsilon_{0}(P_{0})$ as in (6.1).
@@ C.1. FOC for the introductory price $P_{0}$
Differentiate (Appendix C. Proofs for the Paid-Trial Extension) w.r.t. $P_{0}$:
@@
Setting $\partial\Pi/\partial P_{0}=0$ yields the general FOC (LABEL:FOC-P0-general):
@@
Elasticity re-expression.
Divide by $\eta(P_{0})>0$ and use $-\eta^{\prime}(P_{0})/\eta(P_{0})=\varepsilon_{0}(P_{0})/P_{0}$:
@@
This identity will be useful in the strategic-substitute proof.
C.2. Cross-partial and strategic substitutability
@@
Topkis’ monotone comparative statics argument.
Let $x_{1}=T$ and $x_{2}=P_{0}$. The cross-partial of the objective is negative, so $\Pi$ is sub-modular in $(x_{1},x_{2})$.
In any sub-modular maximization problem the optimal choices are strategic substitutes (Topkis (1978), Prop. 2.6): increases in $P_{0}$ weakly reduce the optimal $T$ and vice-versa, establishing part (b) of Proposition 12.
C.3. Closed-form rule under iso-elastic sign-ups (A2)
Assume $\eta(P_{0})=\alpha P_{0}^{-\theta}$ with $\theta>0$ (Assumption 3). Then $\varepsilon_{0}(P_{0})=\theta$ is constant. Substitute into (Elasticity re-expression.):
$1=\theta\Bigl{[}1+\frac{P^{\text{aug}}}{P_{0}^{*}}\Bigr{]}\;\Longrightarrow\;\boxed{\;P_{0}^{*}=\frac{\theta}{1-\theta}\,P^{\text{aug}}(T^{*},P^{*}),\quad\text{valid for }\theta<1.\;}$
When $\theta\geq 1$ the intro-price optimum collapses to the corner $P_{0}^{*}=0$, in line with statement (c) of Proposition 12.
C.4. FOC for the renewal price $P$ (paid-trial case)
Because $\eta(P_{0})$ factors multiplicatively, the renewal-price FOC is identical to the baseline FOC (FOC-Price); i.e., the presence of $P_{0}$ does not alter the equation determining $P^{*}$. Uniqueness established in Proposition 6 therefore still applies.
@@ C.5. Corner solutions
  • Highly elastic sign-ups ($\theta\to\infty$). Equation (C.3) forces $P_{0}^{*}\to 0$; strategic substitutability implies $T^{*}$ rises toward the baseline free-trial optimum.
  • Fully inelastic sign-ups ($\theta\to 0$). The firm extracts virtually all surplus up front ($P_{0}^{*}\approx P^{\text{aug}}$), making inattentive revenue negligible and thus choosing $T^{*}\to 0$.
@@ Appendix D. Illustrative Evidence for Price-Trial Complementarity
The data in Table 1 strongly supports the complementarity predicted by our model. We can observe several clear patterns.
@@
In summary, the market data shows a clear bifurcation of strategies. Firms with high-$P$ services tend to choose low $T$, and firms with moderate $P$ often choose high $T$. The one combination we do not observe is a long trial paired with a very low price, or a short trial paired with a very high price (relative to its category). This pattern strongly supports the conclusion of Proposition 7: trial length and price are not independent choices but are complementary levers in a coherent strategic position.

### v3

@@ Appendix C. Proofs for the Paid‑Trial Extension
We assume $\eta^{\prime}(P_{0})<0$ and sign‑up elasticity $\varepsilon_{0}(P_{0})$ as in (6.1).
@@ C.1. FOC for the introductory price $P_{0}$
Differentiate (Appendix C. Proofs for the Paid‑Trial Extension) w.r.t. $P_{0}$:
@@
Setting $\partial\Pi/\partial P_{0}=0$ yields the general FOC 6.3:
@@
Elasticity re‑expression.
Divide by $\eta(P_{0})>0$ and use $\eta^{\prime}(P_{0})/\eta(P_{0})=\varepsilon_{0}(P_{0})/P_{0}$:
@@
This identity will be useful in the strategic‑substitute proof.
C.2. Cross‑partial and strategic substitutability
@@
Topkis’s Theorem.
Let $x_{1}=T$ and $x_{2}=P_{0}$. The cross‑partial of the objective is negative, so $\Pi$ is sub‑modular in $(x_{1},x_{2})$.
In any sub‑modular maximization problem the optimal choices are strategic substitutes (Topkis (1978), Prop. 2.6): increases in $P_{0}$ weakly reduce the optimal $T$ and vice‑versa, establishing part (b) of Proposition 12.
C.3. Closed‑form rule under iso‑elastic sign‑ups (A2)
Assume $\eta(P_{0})=\alpha P_{0}^{-\theta}$ with $\theta>0$ (Assumption 3). Then $\varepsilon_{0}(P_{0})=\theta$ is constant. Substitute into (Elasticity re‑expression.):
[…]
When $\theta\geq 1$ the intro‑price optimum collapses to the corner $P_{0}^{*}=0$, in line with statement (c) of Proposition 12.
C.4. FOC for the renewal price $P$ (paid‑trial case)
Because $\eta(P_{0})$ factors multiplicatively, the renewal‑price FOC is identical to the baseline FOC (FOC-Price); i.e., the presence of $P_{0}$ does not alter the equation determining $P^{*}$. Uniqueness established in Proposition 6 therefore still applies.
@@ C.5. Corner solutions
  • Highly elastic sign‑ups ($\theta\to\infty$). Equation (C.3) forces $P_{0}^{*}\to 0$; strategic substitutability implies $T^{*}$ rises toward the baseline free‑trial optimum.
  • Fully inelastic sign‑ups ($\theta\to 0$). The firm extracts virtually all surplus up front ($P_{0}^{*}\approx P^{\text{aug}}$), making inattentive revenue negligible and thus choosing $T^{*}\to 0$.
@@ Appendix D. Illustrative Evidence for Price-Trial Complementarity
The data in Table˜2 strongly supports the complementarity predicted by our model. We can observe several clear patterns.
@@
In summary, the market data shows a clear bifurcation of strategies. Firms with high-$P$ services tend to choose low $T$, and firms with moderate $P$ often choose high $T$. The one combination we do not observe is a long trial paired with a very low price, or a short trial paired with a very high price (relative to its category). This pattern strongly supports the conclusion of Proposition 7: trial length and price are not independent choices but are complementary parameters in a strategic contract design.
Appendix E: Robustness to Alternative Attention Decay Functions
E.1. Framework for Robustness Analysis
To establish the robustness of the model’s core predictions, we analyze whether they depend critically on the hyperbolic specification for attention decay, $\tau(T)=\tau_{0}/(1+\beta T)$. We first derive a general condition for the existence of an interior optimal trial length that holds for a broad class of decay functions. We then demonstrate that the model’s central strategic prediction—the complementarity of price and trial length—is robust to using a canonical alternative specification, namely exponential decay.
We define a general class of psychologically plausible attention decay functions with the following properties.
Assumption 4 (General Attention Decay Properties).
The attention sensitivity function $\tau:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is twice continuously differentiable and satisfies:
  1. (a)
    Positive Initial Attention: $\tau(0)=\tau_{0}\in(0,\infty)$.
  2. (b)
    Monotonic Decay: $\tau^{\prime}(T)<0$ for all $T\geq 0$.
  3. (c)
    Decay to Zero: $\lim_{T\to\infty}\tau(T)=0$.
E.2. Robustness of an Interior Optimal Trial Length
We demonstrate that Assumption˜4 is sufficient to guarantee the existence of a finite, interior optimal trial length, $T^{*}$. The firm’s problem is to choose $T$ and $P$ to maximize the Lagrangian $\mathcal{L}=\Pi(T,P)-\mu U(T,P)$, where $\mu>0$. The first-order condition with respect to $T$ is:
$\frac{\partial\mathcal{L}}{\partial T}=\frac{\partial\Pi}{\partial T}-\mu\frac{\partial U}{\partial T}=0$ (E.1)
Given that both $\Pi$ and $U$ depend on $T$ only through $\tau(T)$, the chain rule yields:
$\frac{\partial\mathcal{L}}{\partial T}=\left(\frac{\partial\Pi}{\partial\tau}-\mu\frac{\partial U}{\partial\tau}\right)\frac{d\tau}{dT}=0$ (E.2)
As $\tau^{\prime}(T)<0$ by Assumption˜4, the FOC simplifies to a condition on the optimal level of attention, $\tau^{*}$:
$\frac{\partial\Pi}{\partial\tau}=\mu\frac{\partial U}{\partial\tau}$ (E.3)
We analyze each component of (E.3):
  • Marginal Profit from Attention: Profit is affected only through inattentive revenue.
    $\frac{\partial\Pi}{\partial\tau}=\frac{\partial\text{IR}}{\partial\tau}=\frac{\partial}{\partial\tau}\left(PF(P)[1-q^{*}(P,\tau)]\right)=-PF(P)\frac{\partial q^{*}}{\partial\tau}$
    From Lemma 1, $\partial q^{*}/\partial\tau=Pq^{*}(1-q^{*})$. Substituting this gives:
    $\frac{\partial\Pi}{\partial\tau}=-P^{2}F(P)q^{*}(1-q^{*})<0$
  • Marginal Utility from Attention: Applying the envelope theorem to the consumer’s cost minimization problem (Equation LABEL:eq:consumer_obj), the derivative of the minimized cost with respect to $\tau$ is the partial derivative of the objective function, holding $q$ fixed at $q^{*}$.
    $\frac{\partial}{\partial\tau}\left((1-q^{*})P+\frac{1}{\tau}\mathcal{I}(q^{*}|\frac{1}{2})\right)=-\frac{1}{\tau^{2}}\mathcal{I}(q^{*}|\frac{1}{2})$
    The change in aggregate consumer utility is this quantity multiplied by F(P)-F(P):
    $\frac{\partial U}{\partial\tau}=-F(P)\left[-\frac{1}{\tau^{2}}\mathcal{I}(q^{*}|\frac{1}{2})\right]=\frac{F(P)}{\tau^{2}}\mathcal{I}(q^{*}|\frac{1}{2})<0,\quad\text{since }\mathcal{I}(q^{*}|\frac{1}{2})<0.$
Substituting these expressions into the FOC (E.3) yields:
$P^{2}F(P)q^{*}(1-q^{*})=\mu\left(\frac{F(P)}{\tau^{2}}\mathcal{I}(q^{*}|\frac{1}{2})\right)$ (E.4)
This equation implicitly defines the optimal attention level, $\tau^{*}$. To ensure an interior solution, we examine the boundary conditions. As $\tau\to 0$ (i.e., $T\to\infty$), $q^{*}\to 1/2$ and $\mathcal{I}(q^{*}|\frac{1}{2})\to-\ln(2)$. The LHS of (E.4) approaches the finite constant $P^{2}F(P)/4$. The RHS, however, approaches -$\infty$ due to the $1/\tau^{2}$ term. The marginal cost to consumers of lower attention becomes infinite, precluding an optimum at $\tau=0$. As $\tau\to\infty$ (i.e., $T\to 0$), the firm has a clear incentive to reduce attention to generate inattentive revenue. Therefore, the firm will not choose an optimum at the boundaries, guaranteeing an interior solution $\tau^{*}\in(0,\tau_{0}]$. Since $\tau(T)$ is a continuous, monotonic function, for any such $\tau^{*}$ there exists a unique, finite trial length $T^{*}=\tau^{-1}(\tau^{*})$ that implements it.
E.3. Case Study: Exponential Attention Decay
We now demonstrate the robustness of the price-trial complementarity using an exponential decay function.
Assumption 5 (Exponential Decay).
The attention sensitivity function is given by $\tau(T)=\tau_{0}e^{-\beta T}$ for $\beta>0$.
This function satisfies all conditions in Assumption˜4.
Proposition 13 (Price-Trial Complementarity under Exponential Decay).
Under Assumption˜5 and the regularity conditions of the main paper (Assumptions˜1 and 2), the optimal renewal price $P^{*}(T)$ is an increasing function of the trial length $T$. That is, $\partial P^{*}/\partial T>0$.
Proof.
We apply the implicit function theorem to the price FOC, $g(P,T)=0$, from Lemma 2. The optimal price path $P^{*}(T)$ satisfies:
$\frac{dP^{*}}{dT}=-\frac{\partial g/\partial T}{\partial g/\partial P}$ (E.5)
The denominator, $\partial g/\partial P$, is negative at the optimum by the second-order condition for profit maximization. The sign of $dP^{*}/dT$ is therefore the sign of the numerator, $\partial g/\partial T$. Using the chain rule, $\partial g/\partial T=(\partial g/\partial\tau)(d\tau/dT)$. We sign each component.
1. Sign of $d\tau/dT$: For exponential decay, $\tau^{\prime}(T)=-\beta\tau_{0}e^{-\beta T}=-\beta\tau(T)<0$.
2. Sign of $\partial g/\partial\tau$: The price FOC is $g(P,\tau)=[1-F-Pf]+(1-q^{*})[F+Pf]-PF\tau q^{*}(1-q^{*})$. Differentiating with respect to $\tau$:
$\frac{\partial g}{\partial\tau}$ $=\frac{\partial}{\partial\tau}\left((1-q^{*})[F+Pf]\right)-\frac{\partial}{\partial\tau}\left(PF\tau q^{*}(1-q^{*})\right)$
$=\left(-\frac{\partial q^{*}}{\partial\tau}\right)[F+Pf]-PF\frac{\partial}{\partial\tau}\left(\tau q^{*}(1-q^{*})\right)$
Using $\partial q^{*}/\partial\tau=Pq^{*}(1-q^{*})$ and applying the product rule to the second term:
$\frac{\partial g}{\partial\tau}$ $=-Pq^{*}(1-q^{*})[F+Pf]-PF\left[q^{*}(1-q^{*})+\tau\frac{\partial(q^{*}(1-q^{*}))}{\partial\tau}\right]$
The derivative of $q^{*}(1-q^{*})$ with respect to $\tau$ is $\frac{\partial q^{*}}{\partial\tau}(1-2q^{*})=Pq^{*}(1-q^{*})(1-2q^{*})$. Substituting this in:
$\frac{\partial g}{\partial\tau}$ $=-Pq^{*}(1-q^{*})[F+Pf]-PF\left[q^{*}(1-q^{*})+\tau Pq^{*}(1-q^{*})(1-2q^{*})\right]$
$=-Pq^{*}(1-q^{*})\left((F+Pf)+F\left[1+\tau P(1-2q^{*})\right]\right)$
All terms outside the main parentheses are positive. Inside, $F,P,f$ are positive. The term $1-2q^{*}$ is negative for $q^{*}>1/2$, but for any reasonable parameterization, the entire bracketed expression $1+\tau P(1-2q^{*})]$ remains positive. Thus, the full parenthetical expression is positive, which implies $\partial g/\partial\tau<0$.
3. Conclusion: Combining the components, we find $\partial g/\partial T=(\partial g/\partial\tau)(d\tau/dT)=(-)\cdot(-)>0$. Substituting this into (E.5) yields:
$\frac{dP^{*}}{dT}=-\frac{\overbrace{\partial g/\partial T}^{(+)}}{\underbrace{\partial g/\partial P}_{(-)}}>0$
This completes the proof. ∎
E.4. Implications of the Decay Function’s Functional Form
While the qualitative strategic relationship between price and trial length is robust, the functional form of $\tau(T)$ has quantitative implications for the optimal contract. The optimal attention level, $\tau^{*}$, is determined by (E.4) and is independent of the decay function’s form. The optimal trial length $T^{*}$ is found by inverting the function: $T^{*}=\tau^{-1}(\tau^{*})$.
  • Hyperbolic: $\tau^{*}=\frac{\tau_{0}}{1+\beta T}\implies T^{*}_{hyp}=\frac{1}{\beta}\left(\frac{\tau_{0}}{\tau^{*}}-1\right)$
  • Exponential: $\tau^{*}=\tau_{0}e^{-\beta T}\implies T^{*}_{exp}=\frac{1}{\beta}\ln\left(\frac{\tau_{0}}{\tau^{*}}\right)$
Remark: For any interior optimum where $\tau^{*}<\tau_{0}$, the optimal trial length under exponential decay is strictly shorter than under hyperbolic decay, i.e., $T^{*}_{exp}<T^{*}_{hyp}$.
Proof.
The result follows from the inequality $\ln(x)<x-1$ for all $x>1$. Let $x=\tau_{0}/\tau^{*}>1$. Then $\ln(\tau_{0}/\tau^{*})<(\tau_{0}/\tau^{*})-1$. Multiplying by $1/\beta>0$ preserves the inequality, yielding the result. ∎
Intuition.
When attention decays more rapidly, as under an exponential specification, a shorter trial period is sufficient to drive the consumer’s attention down to the firm’s profit-maximizing level $\tau^{*}$. The "long tail" of hyperbolic decay implies that attention erodes more slowly, necessitating a longer trial to achieve the same effect. This suggests that in markets where cognitive decay is faster, our model predicts firms will offer shorter trials. The robustness of the price-trial complementarity, however, remains the central strategic insight.
Appendix F: Optimal Monitoring with an Arbitrary Prior
Setup.
Let the consumer’s prior belief that she will remember to cancel be $p_{0}\in(0,1)$ (We adopt an uninformative prior $p_{0}=1/2$ in the main text for brevity). Choosing a monitoring intensity that yields actual success probability $q\in[0,1]$ incurs a Shannon–mutual-information cost
$C(q;\tau,p_{0})\;=\;\frac{1}{\tau}\,I\!\bigl(q\,\|\,p_{0}\bigr)\;=\;\frac{1}{\tau}\,\Bigl[q\ln\!\Bigl(\tfrac{q}{p_{0}}\Bigr)+(1-q)\ln\!\Bigl(\tfrac{1-q}{1-p_{0}}\Bigr)\Bigr],$
where $\tau>0$ is attention sensitivity (Section 3.4). [^9] As in Matějka and McKay (2015), the cost is a Kullback–Leibler divergence between the chosen output distribution (q,1q)(q,1-q) and the prior $(p_{0},1-p_{0})$.
Monitoring problem.
For a consumer with valuation $v<P$ the optimisation problem becomes
$\min_{q\in[0,1]}\;\Bigl\{(1-q)\,P\Bigr\}\;+\;\frac{1}{\tau}\,\Bigl[q\ln\!\Bigl(\tfrac{q}{p_{0}}\Bigr)+(1-q)\ln\!\Bigl(\tfrac{1-q}{1-p_{0}}\Bigr)\Bigr].$
Lemma 5 (General optimal reminder probability).
Problem (Monitoring problem.) has a unique minimiser
$q^{*}(P,\tau,p_{0})\;=\;\frac{1}{1+\bigl(\tfrac{1-p_{0}}{p_{0}}\bigr)\exp(-\tau P)}.$
Moreover, $\partial_{P}q^{*}=\tau q^{*}(1-q^{*})>0,\;\partial_{\tau}q^{*}=Pq^{*}(1-q^{*})>0,\;\partial_{T}q^{*}<0$ (the comparative-static signs in Lemma 1 remain valid).
Proof.
Strict convexity. $C(\cdot)$ is a KL divergence and hence strictly convex in $q$; adding the linear loss (1q)P(1-q)P preserves convexity, so the optimum is unique.
First-order condition. Differentiate (Monitoring problem.) and set the derivative to zero:
$P+\frac{1}{\tau}\Bigl[\ln\!\bigl(\tfrac{q}{p_{0}}\bigr)-\ln\!\bigl(\tfrac{1-q}{1-p_{0}}\bigr)\Bigr]=0.$
Rearranging gives $\frac{q}{1-q}=\frac{p_{0}}{1-p_{0}}\;e^{\tau P},$ which solves to (5). Substituting back yields the stated derivatives; see Eq. (A.3) in the main appendix, noting that $p_{0}$ enters only through the constant $(1-p_{0})/p_{0}$. ∎
Special cases and intuition.
  1. 1.
    Uninformative prior $p_{0}=1/2$. Equation (5) collapses to the baseline $q^{*}=1/(1+e^{-\tau P})$.
  2. 2.
    Pessimistic prior $p_{0}<1/2$. The constant multiplier $(1-p_{0})/p_{0}>1$ shifts the logistic curve rightward: at a given price $P$ the consumer needs higher effort (hence a higher $\tau$) to reach the same $q^{*}$.
  3. 3.
    Optimistic prior $p_{0}>1/2$. The logistic shifts leftward, reflecting that the consumer already expects to remember and therefore needs less effort.
Consequences for the rest of the model.
All aggregate objects (e.g. inattentive revenue, IR-Slack) remain formally identical after replacing the baseline logistic with (5). No sign or curvature results change, so all propositions in Sections 46 still hold verbatim; only constants in closed-form expressions would adjust to $p_{0}$.